When 58 J of heat are added to 14 g of a liquid, its temperature rises from 10.4 °C to 13.5 °C.

what is the heat capacity of the liquid
Answer #1

Q = m* S * del T  

Q = heat

S = specific heat

M = mass

Del T = change in temp = 13.5 - 10.4 = 3.1 degree C

Using formula ,

58 J = 14 g * S* 3.1 °c

Heat capacity ( S) = 1.34 J/g °C